Assume that interest rate parity holds. In the spot market 1 Japanese yen = $0.011, while in the 180-day forward market 1 Japanese yen = $0.0118. 180-day risk-free securities yield 1.4% in Japan. What is the yield on 180-day risk-free securities in the United States? Do not round intermediate calculations. Round your answer to two decimal places.
Answer #1
As per IRPT, forward rate = Spot rate*(1+domestic interest rate)/(1+foreign interest rate)
Substituting available values, we have
0.0118 = 0.011*(1+US interest rate)/(1+0.014)
Here, since the quote is for 1 Yen in terms of $,
US is the domestic currency and Yen the foreign
currency.
From the above equation,
US interest rate = 0.0118*1.014/0.011-1 = 8.77%
VERIFICATION:
Forward rate = 0.011*1.0877/1.014 = 0.0118